# One Perturbation is Enough: On Generating Universal Adversarial Perturbations against Vision-Language Pre-training Models

### v3

@@ 4.3 Evaluation on More Downstream Tasks
Image captioning. The objective of IC is to generate text descriptions relevant to the semantic content based on the given image. We use ALBEF, TCL, and BLIP as source models and attack the commonly used captioning model BLIP. Similar to SGA [29], several typical evaluation metrics of IC are calculated to measure the quality of generated captions, including BLEU [35], METEOR [5], ROUGE [26], CIDEr [39], and SPICE [1]. The results in Table 3 demonstrate that our algorithm again displays prominent attack effectiveness, e.g., the crated UAP induces notable drops of 10.2% and 9% in the B@4 and ROUGE_L respectively when transferred from TCL to BLIP.
@@ 4.4 Ablation Study
This part employs ALBEF [23] as the surrogate model and provides ablation studies on Flickr30K. We begin our analysis on the contribution of each proposed technique. Next, we examine the sensitivity of certain hyperparameters.
@@ Appendix B More Training Details
We initialize the noise variable as a matrix. Meanwhile, the initial noise ’s dimensions in the text modality depend on the size of the hidden layer within the specific VLP model. Concretely, we set its dimension to for ALBEF, TCL, BLIP, and X-VLM, while for the CLIP model. When computing the multimodal contrastive loss , the temperature is set as 0.1. The generator is trained over 40 epochs with the Adam optimizer at a learning rate of . Following previous works [29, 41], we employ the attack success rate (ASR) as our quantitative measurement in ITR tasks by computing the extent to which the adversarial perturbations result in victim models’ performance deviations from the clean performance.
@@ Appendix C More Experimental Results
Visual entailment tasks. Given an image and a textual description, visual entailment involves determining whether the textual description can be inferred from the semantic information of the image. We align with previous VLP attacks [48, 41] and conduct experiments on the SNLI-VE [43] dataset using the ALBEF and TCL models. Note that the Baseline represents the clean performance of the target model on the clean data and the orange and green indicate ALBEF and TCL as source models respectively. The results presented in Fig. 6 reveal that C-PGC obtains impressive attack effects by decreasing the average accuracy by nearly 20%.
Notably, [11] has reported a large number of annotation errors in the labels of the SNLI-VE corpus used for VE tasks. Therefore, the presented results are only for experimental integrity and reference purposes.
@@
Performance under Defenses. We next analyze several defense strategies to mitigate the potential harm from C-PGC. Concretely, we totally align with TMM [41] and consider several input preprocessing-based schemes, including image smoothing [10] (Gaussian, medium, average smoothing), JPEG compression [14], NRP [33], and the DiffPure [34], a powerful purification defense using diffusion models. For adversarial text correction, we choose the LanguageTool (LT) [41], which has been widely adopted in various scenarios due to its universality and effectiveness.
The attack results in Table 8 demonstrate that the proposed attack still attains great ASR against different powerful defenses. It also indicates that NRP+LT would be a decent choice to alleviate the threat brought by C-PGC. Another noteworthy finding is that, although DiffPure [34] exhibits remarkable performance in defending attacks in classification tasks, its ability is greatly reduced in V+L scenarios since the denoising process could also diminish some texture or semantic information that is critical for VLP models, thereby acquiring unsatisfactory defense effects.
@@ Appendix E Semantic Similarity Analysis
The basic objective of untargeted adversarial attacks is to fool the victim model to output incorrect predictions [12], while the attacker is supposed to preserve semantic similarity between original and adversarial samples to ensure attack imperceptibility. In our implementation, we follow the rigorous setup in prior works [48, 29, 41] that modify only one single word to preserve the attack stealthiness. To quantitatively analyze the influence, we provide the BERT scores [50], which calculate the P (precision), R (recall), and F1 (F1 score), to measure the semantic distance between 5,000 clean and adversarial sentences in Table 10. Note that we provide existing well-acknowledged sample-specific algorithms Co-Attack [48] and SGA [29] as references.

### v4

@@ 4.3 Evaluation on More Downstream Tasks
Image captioning. The objective of IC is to generate text descriptions relevant to the semantic content based on the given image. We use ALBEF, TCL, and BLIP as source models and attack the commonly used captioning model BLIP. Similar to SGA [lu2023set], several typical evaluation metrics of IC are calculated to measure the quality of generated captions, including BLEU [papineni2002bleu], METEOR [banerjee2005meteor], ROUGE [lin2004rouge], CIDEr [vedantam2015cider], and SPICE [anderson2016spice]. The results in Table 3 demonstrate that our algorithm again displays prominent attack effectiveness, e.g., the crated UAP induces notable drops of 10.2% and 9% in the B@4 and ROUGE_L respectively when transferred from TCL to BLIP.
@@ 4.4 Ablation Study
This part employs ALBEF [li2021align] as the surrogate model and provides ablation studies on Flickr30K. We begin our analysis on the contribution of each proposed technique. Next, we examine the sensitivity of certain hyperparameters.
@@ Appendix B More Training Details
We initialize the noise variable as a matrix. Meanwhile, the initial noise ’s dimensions in the text modality depend on the size of the hidden layer within the specific VLP model. Concretely, we set its dimension to for ALBEF, TCL, BLIP, and X-VLM, while for the CLIP model. When computing the multimodal contrastive loss , the temperature is set as 0.1. The generator is trained over 40 epochs with the Adam optimizer at a learning rate of . Following previous works [lu2023set, wang2024transferable], we employ the attack success rate (ASR) as our quantitative measurement in ITR tasks by computing the extent to which the adversarial perturbations result in victim models’ performance deviations from the clean performance.
@@ Appendix C More Experimental Results
Visual entailment tasks. Given an image and a textual description, visual entailment involves determining whether the textual description can be inferred from the semantic information of the image. We align with previous VLP attacks [zhang2022towards, wang2024transferable] and conduct experiments on the SNLI-VE [xie2019visual] dataset using the ALBEF and TCL models. Note that the Baseline represents the clean performance of the target model on the clean data and the orange and green indicate ALBEF and TCL as source models respectively. The results presented in Fig. 6 reveal that C-PGC obtains impressive attack effects by decreasing the average accuracy by nearly 20%.
Notably, [do2020snli] has reported a large number of annotation errors in the labels of the SNLI-VE corpus used for VE tasks. Therefore, the presented results are only for experimental integrity and reference purposes.
@@
Performance under Defenses. We next analyze several defense strategies to mitigate the potential harm from C-PGC. Concretely, we totally align with TMM [wang2024transferable] and consider several input preprocessing-based schemes, including image smoothing [ding2019advertorch] (Gaussian, medium, average smoothing), JPEG compression [dziugaite2016study], NRP [naseer2020self], and the DiffPure [nie2022diffusion], a powerful purification defense using diffusion models. For adversarial text correction, we choose the LanguageTool (LT) [wang2024transferable], which has been widely adopted in various scenarios due to its universality and effectiveness.
The attack results in Table 8 demonstrate that the proposed attack still attains great ASR against different powerful defenses. It also indicates that NRP+LT would be a decent choice to alleviate the threat brought by C-PGC. Another noteworthy finding is that, although DiffPure [nie2022diffusion] exhibits remarkable performance in defending attacks in classification tasks, its ability is greatly reduced in V+L scenarios since the denoising process could also diminish some texture or semantic information that is critical for VLP models, thereby acquiring unsatisfactory defense effects.
@@ Appendix E Semantic Similarity Analysis
The basic objective of untargeted adversarial attacks is to fool the victim model to output incorrect predictions [dong2018boosting], while the attacker is supposed to preserve semantic similarity between original and adversarial samples to ensure attack imperceptibility. In our implementation, we follow the rigorous setup in prior works [zhang2022towards, lu2023set, wang2024transferable] that modify only one single word to preserve the attack stealthiness. To quantitatively analyze the influence, we provide the BERT scores [zhangbertscore], which calculate the P (precision), R (recall), and F1 (F1 score), to measure the semantic distance between 5,000 clean and adversarial sentences in Table 10. Note that we provide existing well-acknowledged sample-specific algorithms Co-Attack [zhang2022towards] and SGA [lu2023set] as references.